OpenStudy (anonymous):

A 12 ft. ladder leaning against a wall has its base 4 ft. from the wall. Find the sine and the cosine of the angle that the ladder makes with the ground.

OpenStudy (anonymous):

use pathgoren therm

OpenStudy (anonymous):

How should I solve for the sine and cosine after?

OpenStudy (anonymous):

hmm thats the tricky part that i dont know about

OpenStudy (anonymous):

Well, thanks! :)

OpenStudy (ivancsc1996):

Look.\[\sin \theta=4ft/12ft=1/3\]For the cosine you just need to figure out the third side with the pythagoras theorem:\[x=\sqrt{(12ft)^{2}-(4ft)^{2}}=8\sqrt{2}ft\]The you solve for cosine:\[\cos \theta=\frac{ 8\sqrt{2}ft }{ 12ft }=\frac{ 2\sqrt{2} }{ 3 }\approx1\]

OpenStudy (anonymous):

Thank you, @ivancsc1996 ! :)